# Your AI, Not Your View: The Bias of LLMs in Investment Analysis

### v1

@@ Abstract.
In finance, Large Language Models (LLMs) face frequent knowledge conflicts due to discrepancies between pre-trained parametric knowledge and real-time market data. These conflicts become particularly problematic when LLMs are deployed in real-world investment services, where misalignment between a model’s embedded preferences and those of the financial institution can lead to unreliable recommendations. Yet little research has examined what investment views LLMs actually hold. We propose an experimental framework to investigate such conflicts, offering the first quantitative analysis of confirmation bias in LLM-based investment analysis. Using hypothetical scenarios with balanced and imbalanced arguments, we extract models’ latent preferences and measure their persistence. Focusing on sector, size, and momentum, our analysis reveals distinct, model-specific tendencies. In particular, we observe a consistent preference for large-cap stocks and contrarian strategies across most models. These preferences often harden into confirmation bias, with models clinging to initial judgments despite counter-evidence.
Large Language Models, Financial Bias, Knowledge Conflict, Financial Decision-Making, Investment Analysis, Preference, Trustworthy AI
copyright: acmlicensed
@@ 1. Introduction
The rapid advancement of LLMs has spurred a surge of innovation within the financial sector, where they are particularly adept at processing qualitative and unstructured information. Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and aiding in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024). This trend is now evolving towards even greater autonomy through the development of sophisticated LLM-based agents. These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024; Luo et al., 2025; Xiao et al., 2024).
@@
A critical but underexplored issue in financial applications is knowledge conflict. In a domain as fluid and time-sensitive as finance, conflicts between an LLM’s parametric knowledge and real-time market data are frequent. This conflict becomes particularly revealing when the model is presented with a mix of information. Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023). Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its pre-existing knowledge while disregarding the counter evidence.
@@
    What intrinsic preferences do LLMs exhibit towards key financial factors like sector, size, and momentum?
@@
    How do these intrinsic preferences lead to biases when LLMs are forced to make decisions under contradictory evidence?
To answer these questions, this study introduces a three-stage experimental framework designed to systematically elicit and verify LLM biases in investment analysis. In the first stage, we construct targeted arguments for each company, such as positive vs. negative sentiment or momentum vs. contrarian perspectives, to represent competing investment views. In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s latent preferences. In the third stage, we introduce progressively stronger counter evidence to examine the resilience of these preferences, observing how they evolve into rigid, confirmation-biased judgments.
The main contributions of this paper are twofolds. First, we propose a systematic methodology to identify and quantify latent biases in LLMs for financial applications. Second, we provide the first quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent preferences and its stubbornness against contradictory facts. By systematically uncovering these hidden risks, our work lays a critical foundation for developing more transparent and trustworthy financial AI.
@@ 2.1. Knowledge Conflict
A critical vulnerability in LLMs is knowledge conflict, which arises when external, contextual information clashes with the model’s internal, parametric knowledge (Sun et al., 2025; Jin et al., 2024). A significant body of research demonstrates that when faced with such conflicts, LLMs exhibit a strong confirmation bias. Foundational work by (Xie et al., 2023) revealed that LLMs behave as ”stubborn sloths,” clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts. This over-reliance on internal memory is further evidenced by findings that LLMs struggle to suppress their parametric knowledge even when instructed to (Sun et al., 2025) and can exhibit a Dunning-Kruger-like effect, confidently trusting their own faulty beliefs over correct external information (Jin et al., 2024).
@@ 2.2. Financial Biases in LLMs
The presence of these cognitive biases is particularly concerning in the economic and financial domains. Initial research has begun to map their characteristics, with frameworks applying utility theory demonstrating that LLMs are neither perfectly rational nor consistently human-like (Ross et al., 2024). Other studies note that even specialized financial LLMs can exhibit strong irrationalities (Zhou et al., 2024). While this foundational work is critical for establishing the existence of such biases, the methodologies employed often diverge from the complex process of real-world investment analysis. For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by identifying a bias toward recommending specific stocks across thousands of investment scenarios (Zhi et al., 2025).
However, these simplified methods fail to capture the reality of financial analysis, where decisions are made by synthesizing conflicting evidence. A critical gap therefore exists in understanding how an LLM’s preferences behave under such knowledge conflicts. Our work directly addresses this gap by using a more realistic testbed with balanced, contradictory arguments to expose the mechanism by which latent preferences harden into resilient biases.
@@ 3. Methodology
Figure 2. The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting latent preferences through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
[…]
@@ 3.1. Experimental Setup
Accordingly, our investigation is confined to a curated set of 427 prominent stocks, denoted $\mathcal{S}=\{s_{1},s_{2},\dots,s_{427}\}$. These stocks were selected for their continuous listing in the S&P 500 index over the past five years. Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in stored knowledge rather than speculative generation. All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation.
Model Basic Materials Communication Services Consumer Cyclical Consumer Defensive Energy Financial Services Healthcare Industrials Real Estate Technology Utilities
Llama4-Scout (Meta AI, 2024) 0.77 0.90 0.81 0.75 0.96 0.79 0.87 0.90 0.87 0.90 0.89
DeepSeek-V3 (Liu et al., 2024) 0.63 0.81 0.70 0.68 0.82 0.71 0.79 0.79 0.82 0.84 0.79
Qwen3-235B (Yang et al., 2025) 0.46 0.51 0.41 0.49 0.51 0.51 0.50 0.47 0.49 0.61 0.65
Gemini-2.5-flash (Comanici et al., 2025) 0.33 0.46 0.43 0.36 0.54 0.42 0.49 0.49 0.33 0.50 0.39
GPT-4.1 (Achiam et al., 2023) 0.45 0.34 0.41 0.47 0.48 0.42 0.43 0.42 0.42 0.39 0.39
Mistral-24B (Mistral AI, 2024) 0.47 0.32 0.38 0.36 0.39 0.43 0.45 0.41 0.44 0.44 0.44
Table 1. Preference scores ($\pi_{s}$) for each model across various market sectors. For each model, the cell with the lowest preference score is colored red, and the cell with the highest is colored blue. The table highlights that preference strength is a function of model identity. Models such as Llama4-Scout and DeepSeek-V3 show strong and varied preferences, while others like GPT-4.1 and Mistral-24B exhibit a much flatter preference landscape with lower overall scores.
@@ 3.2. Evidence Generation
To construct balanced qualitative and quantitative arguments for each stock $s\in\mathcal{S}$, we leverage Gemini-2.5-Pro (Comanici et al., 2025), a model that is deliberately chosen to be separate from the six LLMs under evaluation. This design ensures neutrality in evidence generation, minimizing alignment with any of the test models. Recent work has highlighted that LLMs can exhibit a strong bias towards LLM-generated content over externally retrieved information (Tan et al., 2024). To neutralize this potential generation bias and ensure that observed preferences are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
Specifically, for every stock $s$, buy evidences ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidences ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidences. To further isolate intrinsic preferences, all evidences are engineered with a uniform linguistic structure and a fixed intensity parameter $I=5\%$. Thus, each buy evidence posits an expected price appreciation of $I$, while each sell evidence anticipates a depreciation of $I$:
$e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,\quad e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-I,$
@@
3.3. Preference Elicitation
This stage aims to elicit the LLM’s latent preferences by leveraging the confirmation bias that emerges during knowledge conflicts. When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its pre-existing parametric knowledge. We deliberately engineer such a conflict using a balanced prompt. The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidences ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity. In this state of informational equilibrium, where external evidence is mutually contradictory, the model’s ultimate decision is hypothesized to be guided by its internal parametric memory regarding the stock $s$. The resulting choice thereby reveals its intrinsic preference.
To quantify this elicited preference, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias. This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the preference score is calculated as:
$\pi_{s}=\left|\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N}\right|,$
where $\pi_{s}\to 1$ indicates a pronounced and consistent preferences.
@@ 3.4. Bias Verification
We aim to verify if a systematically observed group-level preference extends to a consistent bias towards individual stocks within that group. First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average preference.
For any stock $s\in\mathcal{G}^{*}$, evidence that aligns with the group’s established preference (e.g., buy evidence for a buy-preferred group) is termed supporting evidence. Conversely, evidence that opposes this preference is designated as counter evidence. To test if the group preference manifests as a hardened bias, we subject each stock $s\in\mathcal{G}^{*}$ to a test using an imbalanced prompt. This is a prompt where the counter evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial preference. We then measure the decision flip rate, $\phi_{s}$. This verification is conducted from two perspectives: evidence volume and evidence intensity.
Approach 1: Verification by Evidence Volume
One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter evidence than supporting evidence. For example, in a test case for a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s)}|=3$). The flip rate is computed as:
$\phi_{s}^{\text{vol}}=\frac{N_{\text{flip}}^{(s)}}{N},$
where $N_{\text{flip}}^{(s)}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence. A low $\phi_{s}^{\text{vol}}$ signifies stubborn adherence to the bias.
Approach 2: Verification by Evidence Intensity
An alternative approach is to test the model against counter-evidence of a fixed higher intensity while maintaining volumetric parity. For a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
[…]
(Supporting Evidence) $e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,$
(Counter Evidence) $e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-(I+\Delta).$

### v4

@@ Abstract.
In finance, Large Language Models (LLMs) face frequent knowledge conflicts arising from discrepancies between their pre-trained parametric knowledge and real-time market data. These conflicts are especially problematic in real-world investment services, where a model’s inherent biases can misalign with institutional objectives, leading to unreliable recommendations. Despite this risk, the intrinsic investment biases of LLMs remain underexplored. We propose an experimental framework to investigate emergent behaviors in such conflict scenarios, offering a quantitative analysis of bias in LLM-based investment analysis. Using hypothetical scenarios with balanced and imbalanced arguments, we extract the latent biases of models and measure their persistence. Our analysis, centered on sector, size, and momentum, reveals distinct, model-specific biases. Across most models, a tendency to prefer technology stocks, large-cap stocks, and contrarian strategies is observed. These foundational biases often escalate into confirmation bias, causing models to cling to initial judgments even when faced with increasing counter-evidence. A public leaderboard benchmarking bias across a broader set of models is available at https://linqalpha.com/leaderboard.
Large Language Models, Financial Bias, Knowledge Conflict, Trustworthy AI, Decision-Making, Investment Analysis, Preference
journalyear: 2025copyright: ccconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770375isbn: 979-8-4007-2220-2/2025/11ccs: Computing methodologies Natural language processingccs: Applied computing Economics
@@ 1. Introduction
The rapid advancement of LLMs has spurred a surge of innovation within the financial sector, where they are particularly adept at processing qualitative and unstructured information. Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and assisting in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024). This trend is now evolving towards even greater autonomy through the development of sophisticated LLM-based agents. These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024).
@@
A critical but underexplored issue in financial applications is knowledge conflict. In a domain as fluid and time-sensitive as finance, conflicts between an LLM’s parametric knowledge and real-time market data are frequent. This conflict becomes particularly revealing when the model is presented with a mix of information. Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter-evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023). Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its internal knowledge while disregarding the counter-evidence.
@@
    What biases do LLMs exhibit towards key financial factors like sector, size, and momentum?
@@
    What issues arise from these biases when LLMs are forced to make decisions with conflicting evidence?
To answer these questions, this study introduces a three-stage experimental framework designed to systematically elicit and verify LLM biases in investment analysis. In the first stage, we generate equally compelling but opposing arguments for each stock, such as positive vs. negative sentiment or momentum vs. contrarian perspectives, to represent competing investment views. In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s underlying biases. In the third stage, we introduce progressively stronger counter-evidence to examine the problems that arise from these biases, observing how they evolve into rigid, confirmation-biased judgments.
The main contributions of this paper are twofold. First, we propose a systematic methodology to identify and quantify latent biases in LLMs for financial applications. Second, we provide the quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent biases and the critical issues that arise, such as a stubborn refusal to revise judgments despite increasing counter-evidence. This work serves as a critical reminder that for financial AI, trustworthiness is a benchmark as vital as performance. By systematically uncovering these hidden risks, our work lays a critical foundation for developing more transparent and trustworthy financial AI.
@@ 2. Background
This section provides background on knowledge conflict and financial biases in LLMs to contextualize our investigation into their intersection.
@@ 2.1. Knowledge Conflict
A critical vulnerability in LLMs is knowledge conflict, which arises when external, contextual information clashes with the model’s internal, parametric knowledge (Sun et al., 2025; Jin et al., 2024). A significant body of research demonstrates that when faced with such conflicts, LLMs exhibit a strong confirmation bias. Foundational work by (Xie et al., 2023) revealed that LLMs behave as "stubborn sloths," clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts. This over-reliance on internal memory is further evidenced by findings that LLMs struggle to suppress their parametric knowledge even when instructed to (Sun et al., 2025) and can exhibit a Dunning-Kruger-like effect, confidently trusting their own faulty beliefs over correct external information (Jin et al., 2024).
@@ 2.2. Financial Biases in LLMs
The presence of these biases is particularly concerning in the economic and financial domains (Bini et al., 2025; Cao et al., 2025; Chen et al., 2024; Cook and Kazinnik, 2025). Recent evidence suggests that alignment tuning can shift LLMs’ risk preferences, influencing their financial decision-making (Ouyang et al., 2024). Initial research has begun to map their characteristics, with frameworks applying utility theory demonstrating that LLMs are neither perfectly rational nor consistently human-like (Ross et al., 2024). Other studies note that even specialized financial LLMs can exhibit strong irrationalities (Zhou et al., 2024). While this foundational work is critical for establishing the existence of such biases, the methodologies employed often diverge from the complex process of real-world investment analysis. For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by detecting preferential stock recommendations across numerous scenarios (Zhi et al., 2025).
However, existing simplified approaches overlook the core of real-world investment analysis: synthesizing conflicting signals. We address this gap by designing a realistic testbed that presents LLMs with balanced, contradictory arguments. By progressively introducing stronger counter-evidence, we then observe how initial biases lead to flawed decision-making. This approach provides a systematic framework for diagnosing the practical risks posed by LLM biases in finance.
@@ 3. Methodology
This study adopts a three-stage experimental framework, as illustrated in Figure 2, to examine whether biases in LLMs cause skewed financial decisions. All experiments utilize a standardized prompt structure, $\mathcal{P}=(T,C,A)$, comprising three components: a fixed Task ($T$) instructing the model to make an investment decision, a variable Context ($C$) containing the evidence set, and a fixed set of permissible Actions ($A$) defined as $\{\text{buy},\text{sell}\}$. Our methodology is designed to probe the model’s behavior when faced with conflicting information within this framework.
Figure 2. The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting bias through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
@@ 3.1. Experimental Setup
Accordingly, our investigation is confined to a curated set of 427 prominent stocks, denoted $\mathcal{S}=\{s_{1},s_{2},\dots,s_{427}\}$. These stocks were selected for their continuous listing in the S&P 500 index over the past five years. Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in internal knowledge rather than speculative generation. All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation, and each experimental run was conducted three times to ensure the robustness of our results. Detailed specifications of the models used are provided in Appendix A.1.
Figure 3. Sector bias scores for each evaluated LLM. Scores represent the mean of three independent sets of 10 trials; the standard deviation in parentheses reflects the variation across the three sets. Green indicates a positive (buy) bias and red a negative (sell) bias. A strong bias toward the Technology sector is evident in most models.
@@ 3.2. Evidence Generation
To construct balanced qualitative and quantitative arguments for each stock $s\in\mathcal{S}$, we leverage Gemini-2.5-Pro (Comanici et al., 2025), a model that is deliberately chosen to be separate from the six LLMs under evaluation. This design ensures neutrality in evidence generation, minimizing alignment with any of the test models. Recent work has highlighted that LLMs can exhibit a strong bias towards LLM-generated content over externally retrieved information (Tan et al., 2024). To neutralize this potential generation bias and ensure that observed bias are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
Specifically, for every stock $s$, buy evidence ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidence ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidence. To further isolate bias, all evidence is engineered with a uniform linguistic structure and explicitly states a fixed intensity parameter $I=5\%$. Thus, each buy evidence posits an expected price appreciation of $I$, while each sell evidence anticipates a depreciation of $I$:
[…]
@@
3.3. Bias Elicitation
This stage aims to elicit the LLM’s underlying bias by leveraging the confirmation bias that emerges during knowledge conflicts. When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its parametric knowledge. We deliberately engineer such a conflict using a balanced prompt. The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity. In this state of informational equilibrium, where external evidence is mutually contradictory, the model’s ultimate decision is hypothesized to be guided by its internal parametric memory regarding the stock $s$. The resulting choice thereby reveals its inherent bias. See Appendix A.2 for detailed prompt example.
To quantify this elicited bias, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias. This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the bias score($\pi_{s}$) is calculated as:
(2) $\pi_{s}=\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N_{\text{buy}}^{(s)}+N_{\text{sell}}^{(s)}},$
where the resulting score $\pi_{s}$ ranges from -1 to 1. A value approaching 1 indicates a pronounced bias towards buying, while a value approaching -1 indicates a pronounced bias towards selling.
@@ 3.4. Bias Verification
The goal of this stage is to verify the problems that arise from groups exhibiting high bias scores. First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average bias score.
To test what problems this group-level bias causes in the decision-making process, let $s^{*}$ denote any stock from this most-biased group ($s^{*}\in\mathcal{G}^{*}$). Evidence that aligns with the group’s established bias for $s^{*}$ (e.g., buy evidence for a buy-biased group) is termed supporting evidence. Conversely, evidence that opposes this bias is designated as counter-evidence. We then subject each stock $s^{*}\in\mathcal{G}^{*}$ to a test using an imbalanced prompt. This is a prompt where the counter-evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial bias. We then measure the decision flip rate, $\phi_{s^{*}}$. This verification is conducted from two perspectives: evidence volume and evidence intensity.
3.4.1. Approach 1: Verification by Evidence Volume
One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter-evidence than supporting evidence. For example, in a test case for a stock $s^{*}\in\mathcal{G}^{*}$ from a buy-biased group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s^{*})}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s^{*})}|=3$). The flip rate is computed as:
(3) $\phi_{s^{*}}^{\text{vol}}=\frac{N_{\text{flip}}^{(s^{*})}}{N},$
where $N_{\text{flip}}^{(s^{*})}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence. A low $\phi_{s^{*}}^{\text{vol}}$ signifies that the model’s intrinsic bias is strong enough to outweigh the volumetric majority of counter-evidence.
3.4.2. Approach 2: Verification by Evidence Intensity
An alternative approach is to test the model against counter-evidence of a fixed higher intensity while maintaining volumetric parity. For a stock $s^{*}\in\mathcal{G}^{*}$ from a buy-biased group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
This creates asymmetric conflict. For example, for a baseline intensity of $I=5\%$ and an increment of $\Delta=5\%$, the intensified level would be $10\%$. The expectations are:
(4) (Supporting Evidence) $e_{\text{buy},i}^{(s^{*})}:\mathbb{E}[\Delta p^{(s^{*})}]=+I,$
(Counter-Evidence) $e_{\text{sell},i}^{(s^{*})}:\mathbb{E}[\Delta p^{(s^{*})}]=-(I+\Delta).$